40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 17 16:44:04 2015, Koji, Update, LSC, Delay line installed again 
    Reply  Tue Feb 17 19:49:51 2015, Koji, Update, LSC, Delay line un-installed again 
       Reply  Mon Feb 23 21:57:13 2015, Koji, Update, LSC, Delay line installed again (experiment, round 1) 
          Reply  Tue Feb 24 01:57:01 2015, Jenne, Update, LSC, Modelled effect of relative modulation phase ChangeInAngle_PRCLvsMICH.pngREFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.pngREFL33_SensMat_vsModulationPhase.pngREFL33_PRCL_nominalVsLoweredGamma.pngREFL33_MICH_nominalVsLoweredGamma.png
             Reply  Thu Feb 26 00:20:54 2015, ericq, Update, LSC, Modelled effect of relative modulation phase 33budget_canceled.png33budget_nominal.png2015-02-ModPhase.zip
                Reply  Thu Feb 26 01:51:39 2015, ericq, Update, LSC, Sideband HOMs C1_HOMcurves_PR.pngC1_HOMcurves_Y.pngC1_HOMcurves_X.pngC1_HOMlist.zip
                Reply  Thu Feb 26 01:53:35 2015, Koji, Update, LSC, Modelled effect of relative modulation phase 
       Reply  Tue Feb 24 23:39:16 2015, Jenne, Update, LSC, Delay line re-installed again 
          Reply  Wed Feb 25 04:59:47 2015, Jenne, Update, LSC, Delay line re-installed, measurements round 2 SensMat_PRMI_24Feb2015.pdf
             Reply  Wed Feb 25 12:16:27 2015, Koji, Update, LSC, Delay line re-installed, measurements round 2 
                Reply  Wed Feb 25 14:18:28 2015, rana, Update, LSC, Delay line re-installed, measurements round 2 
          Reply  Wed Feb 25 14:51:13 2015, Jenne, Update, LSC, Delay line un-installed again 
             Reply  Wed Feb 25 20:00:39 2015, Jenne, Update, LSC, Delay line un-installed again - sensing matrix comparisons 
                Reply  Wed Feb 25 23:48:57 2015, rana, Update, LSC, some thoughts 
                   Reply  Thu Feb 26 20:02:46 2015, Jenne, Update, LSC, re: some thoughts 
Message ID: 11080     Entry time: Thu Feb 26 20:02:46 2015     In reply to: 11071
Author: Jenne 
Type: Update 
Category: LSC 
Subject: re: some thoughts 

I have clarified my elog from last night to indicate that the sensing matrix in the "33MHz cancellation" configuration was measured with the PRMI held on REFL55 I&Q.

Also, I just re-read my control room notes from yesterday, and I typed the wrong demod phases into the table last night.  The elog has been edited.  Most significantly, the REFL33 demod phase did not change by 70 degrees.  It did change by 10 degrees, but that is likely from the fact that it was formerly tuned for PRCL in PRFPMI, and last night I tuned it for PRCL in PRMI-only.

ELOG V3.1.3-